A large city requires about 1.0 X 10^14 J of electrical enegery per day. if this produced by a nuclear powered generating station that converts 33% of the heat it porduces into electrical enegery, how much mass converted to energy each day?

A 5.0 g

B 3.4 g

C 0.61 g

D 12 g

E 40 g
Answer #1

answer:

option B) m = 3.4 g

explanation:

the energy is related by:

|\Delta E | = \Delta m \left(\frac{931.5\;MeV}{1\;u} \right )

we have that:

|\Delta E | = \frac{1.0\times10^{14}\;J}{0.33} = 3.030\times10^{14}\;J = 1.891\;MeV

thus:

\Delta m = \frac{|\Delta E|}{931.5\;MeV} = \frac{1.891\;MeV}{931.5\;MeV} = 2.030\times10^{24}\;u

convert to grams:

\Delta m = 3.37\;g